find the roots of the quadratic equation by factoring. X^2-9/4x-25
Answer by Alan3354(69443)

find the roots of the quadratic equation by factoring. X^2-9/4x-25
-----------
1st, it's not an equaion, there's no equal sign.
2nd, is it (9/4)x, or 9/(4x) ? or 9/(4x-25) ?
Or (x^2 - 9)/4x ?
Parentheses are free, use some.
---------------------------
Via email: it is x^2-(9/4)x-25
I'll assume it's = 0
x^2 - 2.25x - 25 = 0
4x^2 - 9x - 100 = 0
---------------------
Factoring 4x^2 - 9x - 100 is tricky.
Since the middle term is odd, the 9, it can't be (2x ...)*(2x ...), that would make both x terms even, and they wouldn't add or subtract to an odd number, so that's a clue. That makes it (x ....)*(4x...).
By the same reasoning, one of the factors of 100 has to be odd, either 1, 5 or 25, and the odd number has to be in the (4x...) binomial.
----------------------
(4x - 25)*(x + 4) = 0
x = -4, 6.25